Use the discriminant to determine all values of k which would result in the equation kx^2+2x+1=0 having equal roots
Answer by math_tutor2020(3816)


For the quadratic
ax^2 + bx + c = 0

The discriminant is
D = b^2 - 4ac

In our case,
a = k
b = 2
c = 1

So,
D = b^2 - 4ac
D = 2^2 - 4k*1
D = 4 - 4k

When we have two equal roots, aka repeated root, the discriminant is D = 0
D = 4 - 4k
0 = 4 - 4k
4k = 4
k = 4/4
k = 1

Answer: k = 1